# Adversarial Observability and Performance Tradeoffs in Optimal Control

### v1

@@ Abstract
We develop a feedback controller that minimizes the observability of a set of adversarial sensors of a linear system, while adhering to strict closed-loop performance constraints. We quantify the effectiveness of adversarial sensors using the trace of their observability Gramian and its inverse, capturing both average observability and the least observable state directions of the system. We derive theoretical lower bounds on these metrics under performance constraints, characterizing the fundamental limits of observability reduction as a function of the performance tradeoff. Finally, we show that the performance-constrained optimization of the Gramian’s trace can be formulated as a one-shot semidefinite program, while we address the optimization of its inverse through sequential semidefinite programming. Simulations on an aircraft show how the proposed scheme yields controllers that deteriorate adversarial observability while having near-optimal closed-loop performance.
@@ 1 Introduction
A potential solution is moving target defense – a security method that aims to make the system’s evolution unpredictable and thus disrupts adversarial observability [3, 4, 5, 6]. The underlying concept is that, to impede an adversary from performing system reconnaissance or from predicting future trajectories, one should constantly change the system’s control mode. This induces a randomization in the evolution of the system that renders future tasks of the system difficult to identify. However, switching between different system modes can create discontinuities and transients in control execution, potentially degrading performance and compromising safety. Moreover, when the system has to commit to meeting a single objective, employing different control modes may not always be feasible.
To obstruct the state reconstruction capability of adversarial sensors, several studies have investigated the design of a single controller that reduces the observability of such sensors [7, 8, 9, 10]. The main idea they exploit is a well-known fact in control theory, namely that appropriately chosen state-feedback can destroy the observability of a system from a given output. However, they view the destruction of adversarial observability as a binary and hard constraint, which can result in unspecified loss of closed-loop performance as a tradeoff. Moreover, binary observability constraints do not capture directional structure: by forcing unobservability in some directions, one may inadvertently enhance sensitivity along other, more observable modes. This motivates the need for a feedback controller that enables continuous, tunable tradeoffs between closed-loop performance and adversarial observability.
@@
We consider a setup where the system’s operator, who has full state feedback, wants to design a linear controller to balance two conflicting objectives: i) to minimize the observability of a set of sensors which an adversary might be using to observe the system; and ii) to stabilize the system to the origin with optimal performance. We cast this dual-objective control problem as a constrained optimization of a metric of unobservability of the adversary’s sensors, subject to a tradeoff hard constraint on the distance of closed-loop performance from optimality. In contrast to the aforementioned studies that treat unobservability as a binary property, we capture observability continuously, drawing from the sensor selection literature and leveraging observability Gramians [11, 12, 13, 14, 15].
@@
Gramians are a prime tool for capturing the controllability and observability of linear systems in a continuous manner that goes beyond rank tests. Observability Gramians, in particular, must be inverted to perform least squares state estimation from partial observations [16], and hence their spectrum directly affects the quality of the resulting state estimate. For this reason, there is a rich body of literature that correlates how one should select the output nodes of a system with the underlying observability Gramian of those nodes [11, 12, 13, 14, 15]. Here, we explore this connection in the context of adversarial observability minimization, where the objective is to design the controller (rather than the sensor configuration) to actively reduce (rather than increase) the adversary’s ability to observe the system.
@@
We specifically quantify the observability of adversarial sensors by adopting the trace of the observability Gramian and its inverse. The former quantifies adversarial observability on average, across all observable directions, while the latter is skewed towards the less observable directions. For both of these metrics, we establish theoretical lower bounds on their optimal values as a function of the tradeoff distance from optimal performance. We show that smaller performance tradeoffs generally yield a sharper decrease in the trace of the observability Gramian, whereas larger performance tradeoffs lead to a steeper increase in the trace of the Gramian’s inverse. Finally, we solve the constrained optimization of these metrics i) precisely for the trace of the Gramian, by proving it boils down to the solution of a semidefinite program (SDP); and ii) approximately for the trace of the Gramian inverse, through a sequential SDP algorithm that we obtain by applying the convex-concave procedure.
Structure: We structure the remainder of the paper as follows. Section 2 formulates the problem of continuously optimizing adversarial observability under performance constraints. Section 3 derives lower bounds as a function of the performance tradeoff parameter. Section 4 develops SDP-based algorithmic solutions for minimizing adversarial observability. Section 5 reports simulation results, and finally, Section 6 concludes the paper.
@@ 2.1 Operator and Adversary Models
$K\in\mathbb{R}^{m\times n}$, which does not only optimally stabilize system (1) to the origin, but also makes $x(t)$ difficult to observe from the output $y(t)$. In other words, the operator wants to design a feedback gain $K$ so that $A+BK$ is Hurwitz and nearly optimal, but also so that $(A+BK,~{}C)$ is as unobservable as possible.
@@
Assumption 1 is standard for the operator’s objective of optimally stabilizing (1) to the origin. Assumption 2 is indicative of the information asymmetry between the operator and the adversary, which the operator will use to make the task of observing $x(t)$ from $y(t)$ more difficult for the adversary. Moreover, while the operator might not know exactly which partial states the adversary can measure (as indicated by $C$), they are likely to know a superset of them, and hence Assumption 2 is not restrictive.
@@ 2.2 Operator’s Control Design Problem
Per the problem formulation, the operator considers two objectives when designing its control gain $K$: i) choosing $K$ to stabilize (1) to the origin optimally; and ii) choosing $K$ to make $(A+BK,~{}C)$ as less observable as possible.
We capture the first of the two objectives with the linear-quadratic cost function
@@
To capture the second objective of minimizing the observability of the pair $(A+BK,~{}C)$, we resort to costs relating to observability Gramians. Following [12, 11], two relevant metrics are
@@
where $W\succeq 0$ is the observability Gramian of the closed-loop pair $(A+BK,~{}C)$, uniquely solving the Lyapunov equation
$(A+BK)^{\mathrm{T}}W+W(A+BK)+C^{\mathrm{T}}C=0.$
@@
The metrics (4) relate to the average energy the system releases through the sensors $C$. Specifically, it holds that $\mathbb{E}\left[\int_{0}^{\infty}y^{\mathrm{T}}(t)y(t)\textrm{d}t\right]=\mathbb{E}\left[\int_{0}^{\infty}x^{\mathrm{T}}(t)C^{\mathrm{T}}Cx(t)\textrm{d}t\right]=\mathrm{tr}(WV)=J_{o1}(K)$, whereas $J_{o2}(K)$ is motivated the inequality $\mathrm{tr}(W^{-1}V^{-1})\geq\frac{n^{2}}{\mathrm{tr}(WV)}$ and the fact that it becomes unbounded when the closed loop is not observable.
Notably, to make the pair $(A+BK,~{}C)$ as unobservable as possible, the operator would typically aim to either maximize $\mathrm{tr}(W^{-1}V^{-1})$ or minimize $\mathrm{tr}(WV)$. The difference between these two metrics lies in the aspect of observability they emphasize. The quantity $\mathrm{tr}(W^{-1}V^{-1})$ is better suited for capturing how close the pair $(A+BK,~{}C)$ is to being unobservable, as it becomes unbounded (infinite) if the system is not observable. This makes it particularly effective when the objective is to suppress the most weakly observable modes. On the other hand, $\mathrm{tr}(WV)$ is better suited for quantifying observability on average, across all directions of the state space. It remains finite even when $(A+BK,~{}C)$ is unobservable, and is thus a preferable metric when one does not place much emphasis on complete unobservability, or when a well-defined metric is needed in both observable and unobservable regimes.
In view of the discussion above, to co-optimize the performance of the closed-loop plant matrix $A+BK$ as well as the unobservability of the pair $(A+BK,~{}C)$, we focus on two optimization problems. The first problem minimizes the first unobservability objective in (4), subject to a constraint that limits the degradation in performance, as quantified by the cost (2) and a tradeoff parameter $\lambda>0$. Specifically, this constraint is $J_{s}(K)-J_{s}(K^{\star})\leq\lambda$, or equivalently
@@
Per the preceding discussion and [17], one should select $\epsilon$ to be close to 00.
Both Problem 1 and 2 are optimization problems with bilinear constraints. Problem 2 is further complicated by its nonlinear cost function. We will provide lower bounds to their values that showcase the fundamental limits of the tradeoff between adversarial observability and performance, as it relates to $\lambda$. Moreover, we will derive an algorithm that exactly solves Problem 1 using an equivalent SDP, and an algorithm that approximately solves Problem 2 through sequential SDP.
3 Adversarial Observability and Performance Tradeoffs
In this section, we study the adversarial observability and performance tradeoff in Problems 1 and 2, by providing lower bounds to the values of these problems as a function of the tradeoff parameter $\lambda$. To that end, define the matrices $Z^{\star},~{}S^{\star},~{}U^{\star}\succ 0$ that solve
@@ 3 Adversarial Observability and Performance Tradeoffs
$\qquad\qquad\qquad+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K})\Big{)}$
$=\mathrm{tr}(Z^{\star}(\tilde{K}^{\mathrm{T}}R\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K}))$
@@
Next, denote the value of Problem 1 with respect to $\lambda$ as $J_{1}(\lambda)$, and that of Problem 2 as $J_{2}(\lambda)$. Using Lemma 1, we provide lower bounds to the values of these problems as functions of the observability-performance tradeoff parameter $\lambda$.
@@
    $J_{2}(\lambda)\geq\frac{1}{1-\frac{2\kappa(V)f(\lambda)\left\lVert V\right\rVert\left\lVert B\right\rVert\|U^{\star}\|\mathrm{tr}(W_{\epsilon}^{0})\mathrm{tr}((W^{0}_{\epsilon})^{-1})}{1-2\mathrm{tr}(S^{\star})f(\lambda)\|B\|}}J_{2}(0).$ (12)
@@
Note that $f(\lambda)\sim O(\lambda)$ for large $\lambda$ and $f(\lambda)\sim O(\sqrt{\lambda})$ for small $\lambda$. Looking at (11), this implies that small performance losses $\lambda$ yield steeper tradeoffs in reducing adversarial observability for the cost (11). On the other hand, the lower bound in Problem 2 can become unbounded as $\lambda$ increases, since a larger $\lambda$ allows the Gramian inverse to become singular. As a result, higher values of $\lambda$ can produce sharper improvements in the objective of Problem 2.
@@
Combining Theorem 1 with Proposition 1 yields lower bounds for Problems 1 and 2 across all values of the tradeoff parameter $\lambda$ that quantify the fundamental limits of adversarial observability reduction under a given level of performance degradation.

### v2

# Adversarial Observability and Performance Trade-offs in Optimal Control

@@ Abstract
We develop a feedback controller that minimizes the observability of a set of adversarial sensors of a linear system, while adhering to strict closed-loop performance constraints. We quantify the effectiveness of adversarial sensors using the trace of their observability Gramian and its inverse, capturing both average observability and the least observable state directions of the system. We derive theoretical lower bounds on these metrics under performance constraints, characterizing the fundamental limits of observability reduction as a function of the performance trade-off. Finally, we show that the performance-constrained optimization of the Gramian’s trace can be formulated as a one-shot semidefinite program, while we address the optimization of its inverse through sequential semidefinite programming. Simulations on an aircraft show how the proposed scheme yields controllers that deteriorate adversarial observability while having near-optimal performance.
@@ 1 Introduction
A potential solution is moving target defense – a security method that aims to make the system’s evolution unpredictable and thus disrupts adversarial observability [16, 24, 12, 5]. The underlying concept is that, to impede an adversary from performing system reconnaissance or from predicting future trajectories, one should constantly change the system’s control mode. This induces a randomization in the evolution of the system that renders future tasks of the system difficult to identify. However, switching between different system modes can create discontinuities and transients in control execution, potentially degrading performance and compromising safety. Moreover, when the system has to commit to meeting a single objective, employing different control modes may not always be feasible. These observations highlight the need for observability-reduction mechanisms that do not rely on mode switching, but instead operate within a single, performance-constrained feedback architecture.
To obstruct the state reconstruction capability of adversarial sensors, several studies have investigated the design of a single controller that reduces observability [26, 27, 20, 1]. Related work has also studied the observability radius as a metric of distance to unobservability [4], network design for controllability metrics [2], and active defense strategies that mislead adversarial observers by modifying their available signals [22]. These approaches, however, either treat observability reduction as a binary constraint, focus on network topology rather than controller synthesis, or assume knowledge of a specific adversarial estimator. Moreover, they do not explicitly quantify the trade-off between closed-loop performance and observability reduction. This motivates feedback controllers that enable continuous, tunable trade-offs between performance and adversarial observability.
@@
We consider a setup where the system’s operator, who has full state feedback, wants to design a linear controller to balance two conflicting objectives: i) to minimize the observability of a set of sensors which an adversary might be using to observe the system; and ii) to stabilize the system to the origin with optimal performance. We cast this dual-objective control problem as a constrained optimization of a metric of unobservability of the adversary’s sensors, subject to a trade-off hard constraint on the distance of closed-loop performance from optimality. In contrast to the aforementioned studies that treat unobservability as a binary property, we capture observability continuously, drawing from the sensor selection literature and leveraging observability Gramians [23, 21, 6, 10, 19].
@@
Gramians are a prime tool for capturing the controllability and observability of linear systems in a continuous manner that goes beyond rank tests. Observability Gramians, in particular, must be inverted to perform least squares state estimation from partial observations [7], and hence their spectrum directly affects the quality of the resulting state estimate. For this reason, there is a rich body of literature that correlates how one should select the output nodes of a system with the underlying observability Gramian of those nodes [23, 21, 6, 10, 19]. Here, we explore this connection in the context of adversarial observability minimization. Unlike sensor selection, which chooses sensor configurations to maximize observability, our objective is to design the controller to actively reduce the adversary’s ability to observe the system.
@@
We specifically quantify the observability of adversarial sensors by adopting the trace of the observability Gramian and its inverse. The former quantifies adversarial observability on average, across all observable directions, while the latter is skewed towards the less observable directions. For both of these metrics, we establish theoretical lower bounds on their optimal values as a function of the trade-off distance from optimal performance. We show that smaller performance trade-offs generally yield a sharper decrease in the trace of the observability Gramian, whereas larger performance trade-offs lead to a steeper increase in the trace of the Gramian’s inverse. Finally, we solve the constrained optimization of these metrics i) precisely for the trace of the Gramian, by proving it boils down to the solution of a semidefinite program (SDP); and ii) approximately for the trace of the Gramian inverse, through a sequential SDP algorithm that we obtain by applying the convex-concave procedure.
@@ 2.1 Operator and Adversary Models
$K\in\mathbb{R}^{m\times n}$, which does not only optimally stabilize system (1) to the origin, but also makes $x(t)$ difficult to observe from the output $y(t)$. In other words, the operator wants to design a feedback gain $K$ so that $A+BK$ is Hurwitz and nearly optimal, but also so that (A+BK,C)(A+BK,~C) is as unobservable as possible.
@@
Assumption 1 is standard for the operator’s objective of optimally stabilizing (1) to the origin. Assumption 2 is indicative of the information asymmetry between the operator and the adversary, which the operator will use to make the task of observing $x(t)$ from $y(t)$ more difficult for the adversary. Knowledge of $C$ is realistic when the adversary’s sensing modality is constrained by physics (e.g., radar can only measure position and velocity [25], while optical telescopes capture external brightness but not onboard telemetry [8]). Finally, while the operator might often not know exactly which partial states the adversary can measure (as indicated by $C$), they are likely to know a superset of them, and hence Assumption 2 is not restrictive; we formalize this robustness in Proposition 2.
@@ 2.2 Operator’s Control Design Problem
Per the problem formulation, the operator considers two objectives when designing its control gain $K$: i) choosing $K$ to stabilize (1) to the origin optimally; and ii) choosing $K$ to make (A+BK,C)(A+BK,~C) as less observable as possible.
We capture the first of the two objectives with the cost function
@@
To capture the second objective of minimizing the observability of the pair (A+BK,C)(A+BK,~C), we resort to costs relating to observability Gramians. Following [21, 23], two relevant metrics are
@@
where $W\succeq 0$ is the observability Gramian of the closed-loop pair (A+BK,C)(A+BK,~C), uniquely solving the Lyapunov equation
$(A+BK)^{\mathrm{T}}W+W(A+BK)+C^{\mathrm{T}}C=0.$ (5)
@@
The metrics (4) relate to the average energy the system releases through the sensors $C$. Specifically, $\mathbb{E}[\int_{0}^{\infty}y^{\mathrm{T}}(t)y(t)\textrm{d}t]=\mathbb{E}[\int_{0}^{\infty}x^{\mathrm{T}}(t)C^{\mathrm{T}}Cx(t)\textrm{d}t]=\mathrm{tr}(WV)=J_{o1}(K)$, whereas $J_{o2}(K)$ is motivated the inequality $\mathrm{tr}(W^{-1}V^{-1})\geq\frac{n^{2}}{\mathrm{tr}(WV)}$ and the fact that it becomes unbounded when the closed loop is not observable.
Remark 3.
Beyond the energy interpretation in Remark 2, the observability Gramian also governs estimation conditioning: in finite-horizon least-squares estimation, the sensitivity of the estimation error scales with the Gramian inverse [7], and more generally, directions that are weakly observable correspond to increased reconstruction uncertainty across estimation frameworks [6, 19]. An alternative measure of unobservability is the observability radius [4], which quantifies the smallest perturbation of $(A,C)$ that renders the system unobservable. However, our objective is not merely to certify proximity to rank loss, but to continuously degrade an adversary’s state-reconstruction capability under a hard performance constraint. Gramian-based metrics provide smooth measures of estimation conditioning [7], which makes them the appropriate choice in this setting.
Notably, to make the pair (A+BK,C)(A+BK,~C) as unobservable as possible, the operator would typically aim to either maximize $\mathrm{tr}(W^{-1}V^{-1})$ or minimize $\mathrm{tr}(WV)$. The difference between these two metrics lies in the aspect of observability they emphasize. The quantity $\mathrm{tr}(W^{-1}V^{-1})$ is better suited for capturing how close the pair (A+BK,C)(A+BK,~C) is to being unobservable, as it becomes unbounded (infinite) if the system is not observable. This makes it particularly effective when the objective is to suppress the most weakly observable modes. On the other hand, $\mathrm{tr}(WV)$ is better suited for quantifying observability on average, across all directions of the state space. It remains finite even when (A+BK,C)(A+BK,~C) is unobservable, and is thus a preferable metric when one does not place much emphasis on complete unobservability, or when a well-defined metric is needed in both observable and unobservable regimes.
In view of the discussion above, to co-optimize the performance of the closed-loop plant matrix $A+BK$ as well as the unobservability of the pair (A+BK,C)(A+BK,~C), we focus on two optimization problems. The first problem minimizes the first unobservability objective in (4), subject to a constraint that limits the degradation in performance, as quantified by the cost (2) and a trade-off parameter $\lambda>0$. Specifically, this constraint is $J_{s}(K)-J_{s}(K^{\star})\leq\lambda$, or equivalently
@@
Per [13], one should select $\epsilon$ to be close to 0.
Both Problem 1 and 2 involve bilinear constraints. Problem 2 is further complicated by its nonlinear cost function. We will provide lower bounds to their values that showcase the fundamental limits of the trade-off between adversarial observability and performance, as it relates to $\lambda$. Moreover, we will derive an algorithm that exactly solves Problem 1 using an equivalent SDP, and an algorithm that approximately solves Problem 2 through sequential SDP.
3 Adversarial Observability and Performance Trade-offs
In this section, we study the adversarial observability and performance trade-off in Problems 1 and 2. We specifically provide lower bounds to the values of these problems as a function of the trade-off parameter $\lambda$, and study their robustness to uncertainty in the adversary’s sensing matrix.
3.1 Quantitative Lower Bounds
Define the matrices $Z^{\star},~S^{\star},~U^{\star}\succ 0$ that solve
@@ 3.1 Quantitative Lower Bounds
$)\displaystyle\qquad\qquad\qquad+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K}$$\Big)$
$)\displaystyle=\mathrm{tr}(Z^{\star}(\tilde{K}^{\mathrm{T}}R\tilde{K}+\tilde{K}^{\mathrm{T}}B^{\mathrm{T}}\tilde{P}+\tilde{P}B\tilde{K}$))
@@
Next, denote the value of Problem 1 with respect to $\lambda$ as $J_{1}(\lambda)$, and that of Problem 2 as $J_{2}(\lambda)$. Using Lemma 1, we provide lower bounds to the values of these problems as functions of the observability-performance trade-off parameter $\lambda$.
@@
    $J_{2}(\lambda)\geq\frac{1}{1-\frac{2\kappa(V)f(\lambda)\left\lVert V\right\rVert\left\lVert B\right\rVert\|U^{\star}\|\mathrm{tr}(W_{\epsilon}^{0})\mathrm{tr}((W^{0}_{\epsilon})^{-1})}{1-2\mathrm{tr}(S^{\star})f(\lambda)\|B\|}}J_{2}(0),$ (13)
@@
Note that $f(\lambda)\sim O(\lambda)$ for large $\lambda$ and $f(\lambda)\sim O(\sqrt{\lambda})$ for small $\lambda$. Looking at (12), this implies that small performance losses $\lambda$ yield steeper trade-offs in reducing adversarial observability for the cost (12). On the other hand, the lower bound in Problem 2 can become unbounded as $\lambda$ increases, since a larger $\lambda$ allows the Gramian inverse to become singular. As a result, higher values of $\lambda$ can produce sharper improvements in the objective of Problem 2. An illustration of the function $f(\lambda)$ is provided in Figure 1 for the first example of the simulations in Section 5.
The lower bounds also depend on a number of other parameters. Most notable is $\left\lVert B\right\rVert$, which dictates that larger actuation authority for the operator allows for sharper decrease in adversarial observability. They also depend on $V$, which captures the interplay between observability minimization and initial condition covariance.
@@
Combining Theorem 1 with Proposition 1 yields lower bounds for Problems 1 and 2 across all values of the trade-off parameter $\lambda$ that quantify the fundamental limits of adversarial observability reduction under a given level of performance degradation.
Figure 1: Illustration of the function $f(\lambda)$.
3.2 Robustness to Adversary’s Sensing Uncertainty
Assumption 2 requires knowledge of the adversary’s sensing matrix $C$. This is often realistic in practice, for instance, when the adversary employs radar with known location and measurement geometry to track an aircraft. In some settings, however, the system operator may only know a superset of the adversary’s possible sensing configurations, rather than the exact matrix $C$.
Solving Problems 1 and 2 using such a superset generally yields conservative (and thus suboptimal) solutions. Nevertheless, any controller obtained in this manner provides a certified upper bound on the adversary’s observability metric, as we establish next.
Proposition 2.
Assume that the adversary’s sensing matrix is $\hat{C}\in\mathbb{R}^{\hat{m}\times n}$, where $\hat{m}\leq m$, and where the rows of $\hat{C}$ are a subset of the rows of $C$. Then, $J_{oj}(K;~\hat{C})\leq J_{oj}(K;~C)$, $j=1,2$.
Proof.
Denote as $W_{C}$ and $W_{\hat{C}}$ the observability Gramians corresponding to $C$ and $\hat{C}$. Then, by definition, we have $(A+BK)^{\mathrm{T}}W_{C}+W_{C}(A+BK)+C^{\mathrm{T}}C=0$ and $(A+BK)^{\mathrm{T}}W_{\hat{C}}+W_{\hat{C}}(A+BK)+\hat{C}^{\mathrm{T}}\hat{C}=0$. Since $\hat{C}^{\textrm{T}}\hat{C}=C^{\textrm{T}}S^{\textrm{T}}SC$ for a row selection matrix $S\in\mathbb{R}^{\hat{m}\times m}$, we have $C^{\textrm{T}}C-\hat{C}^{\textrm{T}}\hat{C}=C^{\textrm{T}}(I-S^{\textrm{T}}S)C\succeq 0$. Hence, $W_{\hat{C}}\preceq W_{C}$, and thus $J_{o1}(K;~\hat{C})=\textrm{tr}(W_{\hat{C}}V)\leq\textrm{tr}(W_{C}V)=J_{o1}(K;~C)$. The fact that $J_{o2}(K;~\hat{C})\leq J_{o2}(K;~C)$ follows similarly since negated inversion preserves matrix ordering.
Another possible sensing uncertainty concerns perturbations in the numerical values of $C$, rather than uncertainty in its rows (e.g., $\hat{C}=C+\delta C$ for some small $\delta C$). In such cases, Problem 1 is inherently more robust since its objective depends quadratically on $C$ (see next section), whereas Problem 2 can be more sensitive to perturbations in $C$, particularly when the Gramian approaches singularity due to the inverse-based objective. One could also address such uncertainty in the worst-case sense through a robust SDP formulation, though this remains beyond the scope of this work.